# Fine-Tuning Diffusion Models via Intermediate Distribution Shaping

### v1

@@ Abstract
Diffusion models are widely used for generative tasks across domains. While pre-trained diffusion models effectively capture the training data distribution, it is often desirable to shape these distributions using reward functions to align with downstream applications. Policy gradient methods, such as Proximal Policy Optimization (PPO), are widely used in the context of autoregressive generation. However, the marginal likelihoods required for such methods are intractable for diffusion models, leading to alternative proposals and relaxations. In this context, we unify variants of Rejection sAmpling based Fine-Tuning (RAFT) as GRAFT, and show that this implicitly performs PPO with reshaped rewards. We then introduce P-GRAFT to shape distributions at intermediate noise levels and demonstrate empirically that this can lead to more effective fine-tuning. We mathematically explain this via a bias-variance tradeoff. Motivated by this, we propose inverse noise correction to improve flow models without leveraging explicit rewards. We empirically evaluate our methods on text-to-image(T2I) generation, layout generation, molecule generation and unconditional image generation. Notably, our framework, applied to Stable Diffusion 2, improves over policy gradient methods on popular T2I benchmarks in terms of VQAScore and shows an $8.81\%$ relative improvement over the base model. For unconditional image generation, inverse noise correction improves FID of generated images at lower FLOPs/image.
@@ 1 Introduction
(a) We conceptualize a Generalized Rejection Sampling (GRS) framework which subsumes various rejection sampling strategies including classical rejection sampling from MCMC literature and Best-of-N. We show that GRS samples from the solution to PPO but with a reshaped reward - fine-tuning using GRS, which we term Generalized Rejection sAmpling Fine-Tuning (GRAFT) enables PPO with marginal KL constraint for diffusion models, despite the marginal likelihoods being intractable.
@@ 2 Preliminaries
PPO for Generative Modeling: Following [14], we introduce PPO in our setting: Consider a state space $\mathcal{X}$, a reward function $r:\mathcal{X}\to\mathbb{R}$ and a reference probability measure $\bar{p}$ over $\mathcal{X}$. Let $\mathcal{P}(X)$ be the set of probability measures over $\mathcal{X}$ and $\alpha\in(0,\infty)$. Define $R^{\mathsf{reg}}:\mathcal{P}(\mathcal{X})\to\mathbb{R}$ by $R^{\mathsf{reg}}(p)=\mathbb{E}_{X\sim p}[r(X)]-\alpha\mathsf{KL}(p||\bar{p})$, where $\mathsf{KL}(\cdot\|\cdot)$ is the KL divergence. PPO aims to obtain
$p^{\text{ppo}}={\operatorname*{arg\,sup}}_{p\in\mathcal{P}(\mathcal{X})}R^{\mathsf{reg}}(p)\,.$ (1)
Using the method of Lagrangian Multipliers, we can show that $p^{\text{ppo}}(x)\propto\exp(r(x)/\alpha)\bar{p}(x)$. In generative modeling literature, $\bar{p}$ is often the law of generated samples from a pre-trained model - fine-tuning is done on the model so as to sample from the tilted distribution $p^{\text{ppo}}$.
PPO via Rejection Sampling: Classical rejection sampling from the Monte Carlo literature [15] can be used to sample from $p^{\mathsf{ppo}}$. We note this folklore result in our setting:
@@
Let $r(x)\leq r_{\max}$ for some $r_{\max}$. Given a sample $Y\sim\bar{p}$, we accept it with probability $\mathbb{P}(\mathsf{Accept}|Y)=\exp(\tfrac{r(Y)-r_{\max}}{\alpha})$. Then, conditioned on $\mathsf{Accept}$, $Y$ is a sample from $p^{\mathsf{ppo}}$.
Lemma 2.1 provides a way to obtain exact samples from $p^{\mathsf{ppo}}$. A well known challenge with this method is sample inefficiency - as often in practice, $\alpha$ is small leading to small acceptance probability. Thus, in practice, methods such as Best-of-N (BoN) which always accept a fixed fraction of samples are used. We now introduce Generalized Rejection sAmpling Fine Tuning (GRAFT), a framework to unify such rejection sampling approaches. More specifically, Lemma 3.2 shows that this still leads to PPO, but with reshaped rewards. We then discuss its utility in the context of diffusion models.
@@ 3 GRAFT: Generalized Rejection sAmpling Fine Tuning
The law of accepted samples under GRS (Def 3.1) given by $p(X^{(1)}=x|X^{(1)}\in\mathcal{A})$ is the solution to the following Proximal Policy Optimization problem:
@@
Lemma 3.2 shows that this acceptance function performs PPO with the reshaped reward $\hat{r}$ satisfying: $\frac{\hat{r}(x)}{\alpha}=\log[\sum_{k=0}^{K-1}\binom{M-1}{k}F(r(x))^{M-k-1}(1-F(r(x)))^{k}\big]$.
Preference Rewards: Setting $M=2$ and $K=1$ in the above formulation gives preference rewards, i.e., $X^{(1)}$ is accepted and $X^{(2)}$ is rejected if $r(X^{(1)})>r(X^{(2)})$ (and vice versa). This strategy performs PPO with the reshaped reward $\frac{\hat{r}(x)}{\alpha}=\log F(r(x))$. Since $F$ is an increasing function, the PPO equivalent monotonically reshapes the reward $r(x)$ to $\log F(r(x))$.
@@
Draw $C_{i}\sim\mathsf{Ber}(p_{i})$ without-replacement among the duplicate/similar samples (i.e, they are marginally Bernoulli but are not independent). Thus, exactly one out of the duplicate molecules are selected almost surely. Applying Lemma 3.2, we conclude that this performs PPO with
@@
While specialized versions of Lemma 3.2 are known in the context of AR models [12], the result is particularly useful in the context of diffusion models. Note that given a sample $x$ along with a prompt $y$, the marginal likelihood $\bar{p}(x|y)$ can be easily computed for AR models. For diffusion models, we only have access to conditional likelihoods along the denoising trajectory of the diffusion process whereas $\mathsf{KL}({p}||\bar{p})$ is intractable. That is, if the denoising process is run from $t_{N}$ to $t_{0}$, we have access to $\bar{p}(x_{t_{i}}|x_{t_{i+1}})$. A commonly used relaxation is the trajectory KL, $\mathsf{KL}({p}(X_{0:T})||\bar{p}(X_{0:T}))$, which can be shown as an upper bound on the marginal KL. As discussed in [7], this constraint can lead to the initial value function bias problem since the KL regularization is with respect to the learned reverse process. It becomes necessary to learn an appropriate tilt even at time $T$. In this context, Lemma 3.2 offers a simple yet effective alternative to implicitly achieve marginal KL regularization.
@@ 4 Partial-GRAFT for Diffusion Models
Having established that GRAFT implicitly performs PPO, we now examine methods to further improve the framework. Continuous diffusion models typically start with Gaussian noise $X_{T}$ at time $T$ and denoise it to the output $X_{0}$ via a discretized continuous time SDE. With $N$ denoising steps, the model constructs a denoising trajectory ${X}_{t_{N}}\to\dots X_{t_{i}}\to\dots\to{X}_{t_{0}}$ ($t_{N}=T$ and $t_{0}=0$), denoted by ${X}_{T:0}$. We now consider the effect of shaping the distribution of an intermediate state $X_{t}$. For the rest of the discussion, we reserve $n$ and $N$ to refer to discrete timesteps, and $t$ and $T$ for continuous time. For any $t\in[0,T]$ denote the marginal density of ${X}_{t}$ by $\bar{p}_{t}({x})$.
@@
The key difference is that the reshaped reward now depends on the expected value of the acceptance function given a partially denoised state $X_{t}$. This tilts $\bar{p}_{t}$ instead of $\bar{p}_{0}$. It is straightforward to modify the PPO rewards corresponding to GRS to that of P-GRS. We illustrate this by instantiating Lemma 4.2 for preference rewards, as done with GRS (Lemma 3.2) above.
@@ 7 Conclusion
We establish GRAFT, a framework for provably performing PPO with marginal KL for diffusion models through rejection sampling. We then introduce P-GRAFT, a principled framework for intermediate distribution shaping of diffusion models and provide a mathematical justification for this framework. Both GRAFT and P-GRAFT perform well empirically, outperforming policy gradient methods on the text-to-image generation task. Further, both frameworks also extend seamlessly to discrete-continuous diffusion models. Finally, we introduce Inverse Noise Correction, a strategy to improve flow models even without explicit rewards and demonstrate significant quality gains even with lower FLOPs/image.
@@ Appendix D Proofs
${p}^{\text{ppo}}(x)=Z_{2}\exp\left(\frac{\hat{r}(x)}{\alpha}\right)\bar{p}(x)$ (7)
where $Z_{2}$ is the normalization constant. Comparing (6) and (7), $\bar{p}^{a}=p^{\text{ppo}}$ whenever:
@@ D.2.2 Preference Rewards
${p}^{\text{ppo}}(x)=Z_{2}\exp\left(\frac{\hat{r}(x)}{\alpha}\right)\bar{p}_{t}(x).$ (9)
Comparing (8) and (9), $\bar{p}^{a}_{t}=p^{\text{ppo}}$ whenever:
@@
We will follow the exposition in [67] for our proofs. $q_{t}$ converges to $q_{\infty}$ as $t\to\infty$. By [67, Lemma 2] applied to the forward process, we conclude that:
@@
Notice that $q_{\infty}$ is the density of the standard Gaussian random variable. Therefore, it satisfies the Gaussian Logarithmic Sobolev inequality [68]. Thus, we can apply [67, Theorem 4] to conclude that for every $s\geq 0$, $\mathsf{KL}(q_{s}||q_{\infty})\leq e^{-2s}\mathsf{KL}(q_{0}||q_{\infty})$. Thus,

### v2

@@ Abstract
Diffusion models are widely used for generative tasks across domains. While pre-trained diffusion models effectively capture the training data distribution, it is often desirable to shape these distributions using reward functions to align with downstream applications. Policy gradient methods, such as Proximal Policy Optimization (PPO), are widely used in the context of autoregressive generation. However, the marginal likelihoods required for such methods are intractable for diffusion models, leading to alternative proposals and relaxations. In this context, we unify variants of Rejection sAmpling based Fine-Tuning (RAFT) as GRAFT, and show that this implicitly performs KL regularized reward maximization with reshaped rewards. We then introduce P-GRAFT to shape distributions at intermediate noise levels and demonstrate empirically that this can lead to more effective fine-tuning. We mathematically explain this via a bias-variance tradeoff. Motivated by this, we propose inverse noise correction to improve flow models without leveraging explicit rewards. We empirically evaluate our methods on text-to-image(T2I) generation, layout generation, molecule generation and unconditional image generation. Notably, our framework, applied to Stable Diffusion 2, improves over policy gradient methods on popular T2I benchmarks in terms of VQAScore and shows an $8.81\%$ relative improvement over the base model. For unconditional image generation, inverse noise correction improves FID of generated images at lower FLOPs/image.
@@ 1 Introduction
(a) We conceptualize a Generalized Rejection Sampling (GRS) framework which subsumes various rejection sampling strategies including classical rejection sampling from MCMC literature and Best-of-N. We show that GRS samples from the solution to the KL regularized reward maximization objective, but with a reshaped reward - fine-tuning using GRS, which we term Generalized Rejection sAmpling Fine-Tuning (GRAFT) enables marginal KL constraint for diffusion models, despite the marginal likelihoods being intractable.
@@ 2 Preliminaries
KL Regularized RL for Generative Modeling: Following [62], we introduce the following reward maximization problem in our setting: Consider a state space $\mathcal{X}$, a reward function $r:\mathcal{X}\to\mathbb{R}$ and a reference probability measure $\bar{p}$ over $\mathcal{X}$. Let $\mathcal{P}(X)$ be the set of probability measures over $\mathcal{X}$ and $\alpha\in(0,\infty)$. Define the KL regularized reward maximization objective $R^{\mathsf{reg}}:\mathcal{P}(\mathcal{X})\to\mathbb{R}$ by $R^{\mathsf{reg}}(p)=\mathbb{E}_{X\sim p}[r(X)]-\alpha\mathsf{KL}(p||\bar{p})$, where $\mathsf{KL}(\cdot\|\cdot)$ is the KL divergence. RL algorithms such as PPO are then used to solve for
$p^{\text{RL}}={\operatorname*{arg\,sup}}_{p\in\mathcal{P}(\mathcal{X})}R^{\mathsf{reg}}(p)\,.$ (1)
Using the method of Lagrangian Multipliers, we can show that $p^{\text{RL}}(x)\propto\exp(r(x)/\alpha)\bar{p}(x)$. In generative modeling literature, $\bar{p}$ is often the law of generated samples from a pre-trained model - fine-tuning is done on the model so as to sample from the tilted distribution $p^{\text{RL}}$.
Optimal Distribution via Rejection Sampling: Classical rejection sampling from the Monte Carlo literature [63] can be used to sample from $p^{\mathsf{RL}}$. We note this folklore result in our setting:
@@
Let $r(x)\leq r_{\max}$ for some $r_{\max}$. Given a sample $Y\sim\bar{p}$, we accept it with probability $\mathbb{P}(\mathsf{Accept}|Y)=\exp(\tfrac{r(Y)-r_{\max}}{\alpha})$. Then, conditioned on $\mathsf{Accept}$, $Y$ is a sample from $p^{\mathsf{RL}}$.
Lemma 2.1 provides a way to obtain exact samples from $p^{\mathsf{RL}}$. A well known challenge with this method is sample inefficiency - as often in practice, $\alpha$ is small leading to small acceptance probability. Thus, in practice, methods such as Best-of-N (BoN) which always accept a fixed fraction of samples are used. We now introduce Generalized Rejection sAmpling Fine Tuning (GRAFT), a framework to unify such rejection sampling approaches. More specifically, Lemma 3.2 shows that this still leads to sampling from the solution of the KL regularized reward maximization objective, but with reshaped rewards. We then discuss its utility in the context of diffusion models.
@@ 3 GRAFT: Generalized Rejection sAmpling Fine Tuning
The law of accepted samples under GRS (Def 3.1) given by $p(X^{(1)}=x|X^{(1)}\in\mathcal{A})$ is the solution to the following KL regularized reward maximization problem:
@@
Lemma 3.2 shows that this acceptance function results in the reshaped reward $\hat{r}$ satisfying: $\frac{\hat{r}(x)}{\alpha}=\log[\sum_{k=0}^{K-1}\binom{M-1}{k}F(r(x))^{M-k-1}(1-F(r(x)))^{k}\big]$.
Preference Rewards: Setting $M=2$ and $K=1$ in the above formulation gives preference rewards, i.e., $X^{(1)}$ is accepted and $X^{(2)}$ is rejected if $r(X^{(1)})>r(X^{(2)})$ (and vice versa). Since $F$ is an increasing function, the reward is monotonically reshaped from $r(x)$ to $\log F(r(x))$.
@@
Draw $C_{i}\sim\mathsf{Ber}(p_{i})$ without-replacement among the duplicate/similar samples (i.e, they are marginally Bernoulli but are not independent). Thus, exactly one out of the duplicate molecules are selected almost surely. Applying Lemma 3.2, we conclude that:
@@
While specialized versions of Lemma 3.2 are known in the context of AR models [2], the result is particularly useful in the context of diffusion models. Note that given a sample $x$ along with a prompt $y$, the marginal likelihood $\bar{p}(x|y)$ can be easily computed for AR models. For diffusion models, we only have access to conditional likelihoods along the denoising trajectory of the diffusion process whereas $\mathsf{KL}({p}||\bar{p})$ is intractable. That is, if the denoising process is run from $t_{N}$ to $t_{0}$, we have access to $\bar{p}(x_{t_{i}}|x_{t_{i+1}})$. A commonly used relaxation is the trajectory KL, $\mathsf{KL}({p}(X_{0:T})||\bar{p}(X_{0:T}))$, which can be shown as an upper bound on the marginal KL. As discussed in [15], this constraint can lead to the initial value function bias problem since the KL regularization is with respect to the learned reverse process (as shown in [15]). It becomes necessary to learn an appropriate tilt even at time $T$. In this context, Lemma 3.2 offers a simple yet effective alternative to implicitly achieve marginal KL regularization.
@@ 4 Partial-GRAFT for Diffusion Models
Having established that GRAFT implicitly samples from the KL regularized reward maximization objective, we now examine methods to further improve the framework. Continuous diffusion models typically start with Gaussian noise $X_{T}$ at time $T$ and denoise it to the output $X_{0}$ via a discretized continuous time SDE. With $N$ denoising steps, the model constructs a denoising trajectory ${X}_{t_{N}}\to\dots X_{t_{i}}\to\dots\to{X}_{t_{0}}$ ($t_{N}=T$ and $t_{0}=0$), denoted by ${X}_{T:0}$. We now consider the effect of shaping the distribution of an intermediate state $X_{t}$. For the rest of the discussion, we reserve $n$ and $N$ to refer to discrete timesteps, and $t$ and $T$ for continuous time. For any $t\in[0,T]$ denote the marginal density of ${X}_{t}$ by $\bar{p}_{t}({x})$.
@@
The key difference is that the reshaped reward now depends on the expected value of the acceptance function given a partially denoised state $X_{t}$. This tilts $\bar{p}_{t}$ instead of $\bar{p}_{0}$. It is straightforward to modify the reshaped rewards corresponding to GRS to that of P-GRS. We illustrate this by instantiating Lemma 4.2 for preference rewards, as done with GRS (Lemma 3.2) above.
@@ 7 Conclusion
We establish GRAFT, a framework for provably performing marginal KL regularized reward maximization for diffusion models through rejection sampling.We then introduce P-GRAFT, a principled framework for intermediate distribution shaping of diffusion models and provide a mathematical justification for this framework. Both GRAFT and P-GRAFT perform well empirically, outperforming policy gradient methods on the text-to-image generation task. Further, both frameworks also extend seamlessly to discrete-continuous diffusion models. Finally, we introduce Inverse Noise Correction, a strategy to improve flow models even without explicit rewards and demonstrate significant quality gains even with lower FLOPs/image.
@@ Appendix D Proofs
${p}^{\text{RL}}(x)=Z_{2}\exp\left(\frac{\hat{r}(x)}{\alpha}\right)\bar{p}(x)$ (7)
where $Z_{2}$ is the normalization constant. Comparing (6) and (7), $\bar{p}^{a}=p^{\text{RL}}$ whenever:
@@ D.2.2 Preference Rewards
${p}^{\text{RL}}(x)=Z_{2}\exp\left(\frac{\hat{r}(x)}{\alpha}\right)\bar{p}_{t}(x).$ (9)
Comparing (8) and (9), $\bar{p}^{a}_{t}=p^{\text{RL}}$ whenever:
@@
We will follow the exposition in [65] for our proofs. $q_{t}$ converges to $q_{\infty}$ as $t\to\infty$. By [65, Lemma 2] applied to the forward process, we conclude that:
@@
Notice that $q_{\infty}$ is the density of the standard Gaussian random variable. Therefore, it satisfies the Gaussian Logarithmic Sobolev inequality [23]. Thus, we can apply [65, Theorem 4] to conclude that for every $s\geq 0$, $\mathsf{KL}(q_{s}||q_{\infty})\leq e^{-2s}\mathsf{KL}(q_{0}||q_{\infty})$. Thus,